# Modeling Microenvironment Trajectories on Spatial Transcriptomics with NicheFlow

### v1

@@ 1 Introduction
Uncovering the principles governing tissue organization across space and time remains one of the most fundamental challenges in biology, with profound implications for evolutionary and developmental studies [1, 2]. While individual cells form the basic units of biological systems, they operate not in isolation but as integral parts of spatially organized microenvironments, functionally distinct neighborhoods, or niches, shaped by cell-to-cell interactions and extracellular components [3, 4, 5]. These spatial microenvironments influence crucial biological processes, from tumor progression to immune infiltration and tissue regeneration [6, 7, 8].
Spatial transcriptomics (ST) has transformed our ability to investigate these tissue architectures by providing single-cell resolution mapping of gene expression while preserving spatial context [9, 10, 11, 12]. This technological breakthrough has enabled researchers to examine the molecular underpinnings of tissue organization with unprecedented detail. However, ST provides only static snapshots of inherently dynamic biological systems. Time-resolved spatial analysis extends ST by capturing how gene expression patterns and cellular arrangements evolve across developmental stages or experimental time [13, 14, 15]. This temporal dimension offers critical insights into the development of tissue organization in health and disease [16] that cannot be inferred from static observations alone.
Despite these technological advances, modeling trajectories on time-resolved spatial slides is complicated, as no direct correspondence exists between cells across slides due to the destructive nature of acquisition practices. Moreover, current computational methods fall short in modeling the evolution of tissue organization at the level of cellular microenvironments. Most approaches infer trajectories by modeling single-cell dynamics using velocity-based models [17, 18, 19] or optimal transport between individual cells [20, 21]. While effective at capturing cell evolution, these cell-centric methods fundamentally miss the coordinated evolution of structured niches within tissues.
@@ 2 Related Work
FM and single-cell transcriptomics. We propose a model based on FM, a framework introduced by several seminal works [22, 23, 24]. Specifically, we adopt a variational view of the FM objective, following Eijkelboom et al. [25], but extend it to mixed-factorized distributions for point cloud generation. Our method, NicheFlow, combines FM with OT, a pairing that has proven effective in modeling cellular data [26, 27, 28, 29]. Unlike these models, however, we focus on point clouds of spatially-resolved transcriptomic profiles. Closest to our approach is Wasserstein FM for point cloud generation [30], applied to reconstruct cellular niches. Yet, that work does not address joint generation of spatial coordinates and cellular states, nor OT-based temporal trajectory prediction, both of which are central to our contribution.
Generative models for spatial transcriptomics. Generative models have been key to spatial tasks such as gene expression prediction from histology slides [31, 32], integration with dissociated single-cell data [32], spatial imputation [33, 34], and perturbation [35]. More recently, LUNA [36] demonstrated strong performance in predicting single-cell spatial coordinates using diffusion models [37] conditioned on transcription data. While related, our model addresses the distinct task of inferring niche trajectories, enabling the simultaneous generation of coordinates and cellular states.
Trajectory inference for spatial transcriptomics. Previous work has explored learning trajectories from spatial slides. Pham et al. [38] proposed a graph-based spatiotemporal algorithm for pseudotime inference, while others leveraged tissue-resolved transcriptomics to estimate cell velocity [17, 18, 19]. Closer to our approach, Klein et al. [20] and Bryan et al. [21] use discrete OT to link cells across time and infer the evolution of cell states from spatially-resolved gene expression. Similarly, DeST-OT [39] aligns spatial slides with semi-relaxed OT couplings, preserving transcriptomic and spatial proximity between ancestor and descendant cells, while SpaTrack [19] uses Fused Gromov-Wasserstein OT [40], balancing transcriptomic and spatial differences based on spatial autocorrelation of features. Unlike our mini-batch deep learning model, these methods do not operate on entire microenvironments and rely on exact OT at the single-cell level, resulting in limitations in scalability and generalization.
@@ 3.1 Optimal Transport with FM
FM [23] is a generative model that transforms a source density $p_{0}$ into a target density $p_{1}$. It operates by learning a time-dependent velocity field $u_{t}({\bm{x}})$ for $t\in[0,1]$ and ${\bm{x}}\in\mathbb{R}^{D}$, which generates a probability path $\{p_{t}\}_{t\in[0,1]}$. This path is constructed such that the marginals at time $t=0$ and $t=1$ match the source and target distributions, i.e., $p_{0}$ and $p_{1}$, respectively. The velocity field induces an Ordinary Differential Equation (ODE), whose solution $\phi_{t}({\bm{x}})$ defines a flow map that transports samples from the source to the target distribution.
@@
Lipman et al. [23] show that regressing the conditional field $u_{t}({\bm{x}}\mid{\bm{x}}_{0},{\bm{x}}_{1})$ is equivalent to learning the marginal field $u_{t}({\bm{x}})$ in expectation. Hence, the FM objective becomes the task of learning the velocity along the conditional probability path between any pair of source and target data points. For a linear conditional probability path, the velocity $u_{t}({\bm{x}}\mid{\bm{x}}_{0},{\bm{x}}_{1})$ has a closed form, and the FM loss is:
@@
In practice, the coupling $\pi({\bm{x}}_{0},{\bm{x}}_{1})$ is instantiated using sample pairs drawn from a mini-batch estimate. When one chooses $\pi^{\star}$ as the OT coupling under a squared Euclidean cost between samples from $p_{0}$ and $p_{1}$, FM approximates the dynamic OT map between source and target densities [41, 26]. Thus, the solution samples $({\bm{x}}_{0},{\bm{x}}_{1})\sim\pi^{\star}$ from the joint distribution approximately follow:
@@ 3.2 Generative OT on incomparable source and target spaces
Klein et al. [29] generalize the OT FM formulation to settings where the source and target distributions are defined on incomparable spaces and propose an approach to generative entropic OT using FM. Given a standard normal noise distribution with samples ${\bm{z}}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{D})$, the authors show that the following sampling procedure:
@@ 3.3 Source-conditioned VFM
Since $u_{t}({\bm{x}}\mid{\bm{x}}_{1})$ is tractable [23], one can recast FM as a variational inference problem, following Eijkelboom et al. [25], by introducing a parameterized approximation $q^{\theta}_{t}({\bm{x}}_{1}\mid{\bm{x}},{\bm{x}}_{0})$ to the true posterior $p_{t}({\bm{x}}_{1}\mid{\bm{x}},{\bm{x}}_{0})$. Integrating the expected velocity in Eq.˜5 over $t\in[0,1]$ enables the generation of target points ${\bm{x}}_{1}$ from noise, conditioned on ${\bm{x}}_{0}$.
@@
$u_{t}({\bm{x}})=u_{t}\left({\bm{x}}\mid\mathbb{E}_{p_{t}({\bm{x}}_{1}\mid{\bm{x}},{\bm{x}}_{0})}\left[{\bm{x}}_{1}\right]\right)$ (7)
@@ 4 NicheFlow
Generative model on structured data. Similar to prior work [30, 36], we consider a generative model over structured point cloud data representing cellular microenvironments. This approach implicitly accounts for spatial correlations between cells, in contrast to models that study the evolution of spatial trajectories at the single-cell level [20].
@@ 5.1.1 Training setup
Datasets. We assess model performance across three spatiotemporal datasets: (i) Mouse embryogenesis [20, 8] and (ii) the axolotl brain development [42], two Stereo-seq datasets profiling the spatially-resolved cellular development of a mouse embryo and axolotl brain across three (E9.5, E10.5 and E11.5) and five time points, respectively. We also consider the (iii) mouse brain ageing dataset [43], profiled with MERFISH [44] across twenty time points (see Secs.˜F.1 and F.2).
@@ 5.1.3 Models and results
Baselines. We compare against what we call SPFlow (Single-Point Flow), a standard FM-based model that predicts temporal trajectories across slides at a single-cell level using an MLP-based velocity field. We also consider RPCFlow (Random Point Cloud Flow), which has the same backbone as NicheFlow, but conditions on randomly sampled point clouds instead of radius neighborhoods. Additionally, we include LUNA [36], a diffusion model for spatial reconstruction from dissociated cells. Note that LUNA does not model temporal dynamics and only generates coordinates from noise with their respective biological annotations. Therefore, we use it as a reference for spatial generation accuracy via the 1NN-F1 metric rather than a proper baseline.
Ablations. We assess different training objectives by comparing standard Conditional Flow Matching (CFM) [26] with two variational formulations modeling posteriors over coordinates and features: Gaussian-only (GVFM) and Gaussian-Laplace (GLVFM). The former uses Gaussian posteriors for coordinates and features. The latter uses the factorized formulation in Sec.˜4.3. For the point-cloud-based methods, we use a fixed value of $\lambda=0.1$ and sampled batches of $64$ regions chosen from the ablations in Sec.˜D.5 and D.6. We present results with a single value for $\lambda$ as this experiment compares different models on reconstructing fixed tissue structures over time, for which we enforce spatial preservation (see Sec.˜4.2). However, we also include model comparisons across multiple values of $\lambda$ in Tab.˜3 for completion.
@@ 5.2 Qualitative evaluation and biological analysis
Experimental setup. We train the model as described in Sec.˜5.1.1. Using the mouse embryonic dataset [8], we select specific microenvironments as source niches for which we want to study the trajectory over time. Specifically, we propose two scenarios for the application of NicheFlow depending on the choice of the OT parameter $\lambda$ (see Sec.˜4.2):
@@
Baseline. We compare NicheFlow with the spatiotemporal framework in moscot [20], which models spatial trajectories at the single-cell level. In contrast to NicheFlow, moscot integrates spatial coordinates directly into the OT formulation using a Fused Gromov-Wasserstein cost [45], where the hyperparameter $\alpha$ controls the trade-off between spatial and feature-based distances (see Sec.˜F.8). A high $\alpha$ increases the influence of spatial distances, whereas in our framework this role is played by the hyperparameter $\lambda$ (see Sec.˜4.2). We provide more details on the selection of the hyperparameter $\alpha$ for our experiments in Sec.˜D.2. Notably, moscot relies on exact OT and learns a transition matrix between source and target samples. As such, it is not a generative model over point clouds like NicheFlow. However, given the overlap in downstream tasks, we consider the comparison relevant.
@@ C.1 Theoretical aspects of Variational Flow Matching
Variational Flow Matching (VFM) [25] relies on the observation that one can write the time-resolved marginal vector field $u_{t}({\bm{x}})$ in FM as the expected conditional field $u_{t}({\bm{x}}\mid{\bm{x}}_{1})$ under the posterior $p_{t}({\bm{x}}_{1}\mid{\bm{x}})$ as:
@@
Efficient simulation. Given a parameterized posterior mean $\mu_{t}^{\theta}$, simulating the generative field in Eq.˜15 is efficient under the linearity condition. For example, in the standard FM setting with straight paths [23], the marginal generative field becomes:
@@ C.2 Factorized posterior
Similar to Eijkelboom et al. [25], in our work, we use a fully factorized posterior, where individual dimensions can follow different families of distributions with finite moments (see Sec.˜4.3). Notably, a factorized approximate posterior over ${\bm{x}}_{1}$ is allowed as a choice for $q^{\theta}_{t}$, since the only requirement to simulate $u_{t}({\bm{x}})$ is for $q_{t}^{\theta}({\bm{x}}_{1}\mid{\bm{x}})$ to match the expectation of $p_{t}({\bm{x}}_{1}\mid{\bm{x}})$ over ${\bm{x}}_{1}$, irrespectively of higher moments or correlations between factors.
@@ C.4 Gaussian and Laplacian Hybrid VFM
In line with Eijkelboom et al. [25], using FM with straight paths enables us to efficiently simulate the marginal generating field using the first moment of the posterior distribution. In other words, for a fully factorized posterior, we only need to parameterize a mean predictor. In our setting, the mean prediction is a neural network $\mu_{t}^{\theta}$ as a time-condition function of a noisy microenvironment $\mathcal{M}$ and a source $\mathcal{M}^{0}$ with outputs:
@@ D.8 Structure-aware evaluation
While we refer to our primary evaluation metrics as PSD and SPD, they correspond to the two asymmetric directions of the Chamfer Distance (CD) - a widely used metric in point cloud generation [46]. PSD measures how closely the predicted points adhere to the ground truth structure (fidelity), while SPD captures how well the prediction covers the full extent of the target (coverage). We compute both metrics over collections of microenvironments tiled across the tissue, providing an indirect assessment of global morphological accuracy.
To directly assess the preservation of internal structure within generated microenvironments, we additionally evaluate Gromov-Wasserstein (GW) and Fused Gromov-Wasserstein (FGW) distances [47, 40]. These metrics compare the intra-point pairwise distances in predicted and ground truth point clouds, capturing latent structural relationships. FGW further incorporates transcriptomic similarity, offering a holistic measure of structural and feature-based alignment.

### v2

@@ 1 Introduction
Uncovering the principles governing tissue organization across space and time remains one of the most fundamental challenges in biology, with profound implications for evolutionary and developmental studies [mayr2019exploring, zinner2020design]. While individual cells form the basic units of biological systems, they operate not in isolation but as integral parts of spatially organized microenvironments, functionally distinct neighborhoods, or niches, shaped by cell-to-cell interactions and extracellular components [ren2023spatial, schaar2024nicheformer, liu2025single]. These spatial microenvironments influence crucial biological processes, from tumor progression to immune infiltration and tissue regeneration [moncada2020integrating, larsson2021spatially, chen2022spatiotemporal].
Spatial transcriptomics (ST) has transformed our ability to investigate these tissue architectures by providing single-cell resolution mapping of gene expression while preserving spatial context [staahl2016visualization, rodriques2019slide, eng2019transcriptome, moses2022museum]. This technological breakthrough has enabled researchers to examine the molecular underpinnings of tissue organization with unprecedented detail. However, ST provides only static snapshots of inherently dynamic biological systems. Time-resolved spatial analysis extends ST by capturing how gene expression patterns and cellular arrangements evolve across developmental stages or experimental time [briggs2018dynamics, wagner2018single, ren2023spatiotemporally]. This temporal dimension offers critical insights into the development of tissue organization in health and disease [papanicolaou2022temporal] that cannot be inferred from static observations alone.
Despite these technological advances, modeling trajectories on time-resolved spatial slides is complicated, as no direct correspondence exists between cells across slides due to the destructive nature of acquisition practices. Moreover, current computational methods fall short in modeling the evolution of tissue organization at the level of cellular microenvironments. Most approaches infer trajectories by modeling single-cell dynamics using velocity-based models [abdelaal2024sirv, long2025spvelo, shen2025inferring] or optimal transport between individual cells [klein2025mapping, bryan2025accurate]. While effective at capturing cell evolution, these cell-centric methods fundamentally miss the coordinated evolution of structured niches within tissues.
@@ 2 Related Work
FM and single-cell transcriptomics. We propose a model based on FM, a framework introduced by several seminal works [liu2023flow, lipman2023flow, albergo2023building]. Specifically, we adopt a variational view of the FM objective, following eijkelboom2024variational, but extend it to mixed-factorized distributions for point cloud generation. Our method, NicheFlow, combines FM with OT, a pairing that has proven effective in modeling cellular data [tong2024improving, pmlr-v238-tong24a, atanackovic2025meta, klein2024genot]. Unlike these models, however, we focus on point clouds of spatially-resolved transcriptomic profiles. Closest to our approach is Wasserstein FM for point cloud generation [haviv2024wasserstein], applied to reconstruct cellular niches. Yet, that work does not address joint generation of spatial coordinates and cellular states, nor OT-based temporal trajectory prediction, both of which are central to our contribution.
Generative models for spatial transcriptomics. Generative models have been key to spatial tasks such as gene expression prediction from histology slides [zhu2025diffusion, wan2023integrating], integration with dissociated single-cell data [wan2023integrating], spatial imputation [haviv2025covariance, li2024stdiff], and perturbation [akbarnejad2025mapping]. More recently, LUNA [yu2025tissue] demonstrated strong performance in predicting single-cell spatial coordinates using diffusion models [ho2020denoising] conditioned on transcription data. While related, our model addresses the distinct task of inferring niche trajectories, enabling the simultaneous generation of coordinates and cellular states.
Trajectory inference for spatial transcriptomics. Previous work has explored learning trajectories from spatial slides. pham2023robust proposed a graph-based spatiotemporal algorithm for pseudotime inference, while others leveraged tissue-resolved transcriptomics to estimate cell velocity [abdelaal2024sirv, long2025spvelo, shen2025inferring]. Closer to our approach, klein2025mapping and bryan2025accurate use discrete OT to link cells across time and infer the evolution of cell states from spatially-resolved gene expression. Similarly, DeST-OT [halmos2025dest] aligns spatial slides with semi-relaxed OT couplings, preserving transcriptomic and spatial proximity between ancestor and descendant cells, while SpaTrack [shen2025inferring] uses Fused Gromov-Wasserstein OT [vayer2020fused], balancing transcriptomic and spatial differences based on spatial autocorrelation of features. Unlike our mini-batch deep learning model, these methods do not operate on entire microenvironments and rely on exact OT at the single-cell level, resulting in limitations in scalability and generalization.
@@ 3.1 Optimal Transport with FM
FM [lipman2023flow] is a generative model that transforms a source density $p_{0}$ into a target density $p_{1}$. It operates by learning a time-dependent velocity field $u_{t}({\bm{x}})$ for $t\in[0,1]$ and ${\bm{x}}\in\mathbb{R}^{D}$, which generates a probability path $\{p_{t}\}_{t\in[0,1]}$. This path is constructed such that the marginals at time $t=0$ and $t=1$ match the source and target distributions, i.e., $p_{0}$ and $p_{1}$, respectively. The velocity field induces an Ordinary Differential Equation (ODE), whose solution $\phi_{t}({\bm{x}})$ defines a flow map that transports samples from the source to the target distribution.
@@
lipman2023flow show that regressing the conditional field $u_{t}({\bm{x}}\mid{\bm{x}}_{0},{\bm{x}}_{1})$ is equivalent to learning the marginal field $u_{t}({\bm{x}})$ in expectation. Hence, the FM objective becomes the task of learning the velocity along the conditional probability path between any pair of source and target data points. For a linear conditional probability path, the velocity $u_{t}({\bm{x}}\mid{\bm{x}}_{0},{\bm{x}}_{1})$ has a closed form, and the FM loss is:
@@
In practice, the coupling $\pi({\bm{x}}_{0},{\bm{x}}_{1})$ is instantiated using sample pairs drawn from a mini-batch estimate. When one chooses $\pi^{\star}$ as the OT coupling under a squared Euclidean cost between samples from $p_{0}$ and $p_{1}$, FM approximates the dynamic OT map between source and target densities [pmlr-v202-pooladian23a, tong2024improving]. Thus, the solution samples $({\bm{x}}_{0},{\bm{x}}_{1})\sim\pi^{\star}$ from the joint distribution approximately follow:
@@ 3.2 Generative OT on incomparable source and target spaces
klein2024genot generalize the OT FM formulation to settings where the source and target distributions are defined on incomparable spaces and propose an approach to generative entropic OT using FM. Given a standard normal noise distribution with samples ${\bm{z}}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{D})$, the authors show that the following sampling procedure:
@@ 3.3 Source-conditioned VFM
Since $u_{t}({\bm{x}}\mid{\bm{x}}_{1})$ is tractable [lipman2023flow], one can recast FM as a variational inference problem, following eijkelboom2024variational, by introducing a parameterized approximation $q^{\theta}_{t}({\bm{x}}_{1}\mid{\bm{x}},{\bm{x}}_{0})$ to the true posterior $p_{t}({\bm{x}}_{1}\mid{\bm{x}},{\bm{x}}_{0})$. Integrating the expected velocity in Eq.˜5 over $t\in[0,1]$ enables the generation of target points ${\bm{x}}_{1}$ from noise, conditioned on ${\bm{x}}_{0}$.
@@
$u_{t}({\bm{x}}\mid\bm{x}_{0})=u_{t}\left({\bm{x}}\mid\mathbb{E}_{p_{t}({\bm{x}}_{1}\mid{\bm{x}},{\bm{x}}_{0})}\left[{\bm{x}}_{1}\right]\right)$ (7)
@@ 4 NicheFlow
Generative model on structured data. Similar to prior work [haviv2024wasserstein, yu2025tissue], we consider a generative model over structured point cloud data representing cellular microenvironments. This approach implicitly accounts for spatial correlations between cells, in contrast to models that study the evolution of spatial trajectories at the single-cell level [klein2025mapping].
@@ 5.1.1 Training setup
Datasets. We assess model performance across three spatiotemporal datasets: (i) Mouse embryogenesis [klein2025mapping, chen2022spatiotemporal] and (ii) the axolotl brain development [axolotl], two Stereo-seq datasets profiling the spatially-resolved cellular development of a mouse embryo and axolotl brain across three (E9.5, E10.5 and E11.5) and five time points, respectively. We also consider the (iii) mouse brain ageing dataset [Sun2025aging], profiled with MERFISH [chen2015spatially] across twenty time points (see Secs.˜F.1 and F.2).
@@ 5.1.3 Models and results
Baselines. We compare against what we call SPFlow (Single-Point Flow), a standard FM-based model that predicts temporal trajectories across slides at a single-cell level using an MLP-based velocity field. We also consider RPCFlow (Random Point Cloud Flow), which has the same backbone as NicheFlow, but conditions on randomly sampled point clouds instead of radius neighborhoods. Additionally, we include LUNA [yu2025tissue], a diffusion model for spatial reconstruction from dissociated cells. Note that LUNA does not model temporal dynamics and only generates coordinates from noise with their respective biological annotations. Therefore, we use it as a reference for spatial generation accuracy via the 1NN-F1 metric rather than a proper baseline.
Ablations. We assess different training objectives by comparing standard Conditional Flow Matching (CFM) [tong2024improving] with two variational formulations modeling posteriors over coordinates and features: Gaussian-only (GVFM) and Gaussian-Laplace (GLVFM). The former uses Gaussian posteriors for coordinates and features. The latter uses the factorized formulation in Sec.˜4.3. For the point-cloud-based methods, we use a fixed value of $\lambda=0.1$ and sampled batches of $64$ regions chosen from the ablations in Sec.˜D.5 and D.6. We present results with a single value for $\lambda$ as this experiment compares different models on reconstructing fixed tissue structures over time, for which we enforce spatial preservation (see Sec.˜4.2). However, we also include model comparisons across multiple values of $\lambda$ in Tab.˜3 for completion.
@@ 5.2 Qualitative evaluation and biological analysis
Experimental setup. We train the model as described in Sec.˜5.1.1. Using the mouse embryonic dataset [chen2022spatiotemporal], we select specific microenvironments as source niches for which we want to study the trajectory over time. Specifically, we propose two scenarios for the application of NicheFlow depending on the choice of the OT parameter $\lambda$ (see Sec.˜4.2):
@@
Baseline. We compare NicheFlow with the spatiotemporal framework in moscot [klein2025mapping], which models spatial trajectories at the single-cell level. In contrast to NicheFlow, moscot integrates spatial coordinates directly into the OT formulation using a Fused Gromov-Wasserstein cost [vayer2019optimal], where the hyperparameter $\alpha$ controls the trade-off between spatial and feature-based distances (see Sec.˜F.8). A high $\alpha$ increases the influence of spatial distances, whereas in our framework this role is played by the hyperparameter $\lambda$ (see Sec.˜4.2). We provide more details on the selection of the hyperparameter $\alpha$ for our experiments in Sec.˜D.2. Notably, moscot relies on exact OT and learns a transition matrix between source and target samples. As such, it is not a generative model over point clouds like NicheFlow. However, given the overlap in downstream tasks, we consider the comparison relevant.
@@ C.1 Theoretical aspects of Variational Flow Matching
Variational Flow Matching (VFM) [eijkelboom2024variational] relies on the observation that one can write the time-resolved marginal vector field $u_{t}({\bm{x}})$ in FM as the expected conditional field $u_{t}({\bm{x}}\mid{\bm{x}}_{1})$ under the posterior $p_{t}({\bm{x}}_{1}\mid{\bm{x}})$ as:
@@
Efficient simulation. Given a parameterized posterior mean $\mu_{t}^{\theta}$, simulating the generative field in Eq.˜15 is efficient under the linearity condition. For example, in the standard FM setting with straight paths [lipman2023flow], the marginal generative field becomes:
@@ C.2 Factorized posterior
Similar to eijkelboom2024variational, in our work, we use a fully factorized posterior, where individual dimensions can follow different families of distributions with finite moments (see Sec.˜4.3). Notably, a factorized approximate posterior over ${\bm{x}}_{1}$ is allowed as a choice for $q^{\theta}_{t}$, since the only requirement to simulate $u_{t}({\bm{x}})$ is for $q_{t}^{\theta}({\bm{x}}_{1}\mid{\bm{x}})$ to match the expectation of $p_{t}({\bm{x}}_{1}\mid{\bm{x}})$ over ${\bm{x}}_{1}$, irrespectively of higher moments or correlations between factors.
@@ C.4 Gaussian and Laplacian Hybrid VFM
In line with eijkelboom2024variational, using FM with straight paths enables us to efficiently simulate the marginal generating field using the first moment of the posterior distribution. In other words, for a fully factorized posterior, we only need to parameterize a mean predictor. In our setting, the mean prediction is a neural network $\mu_{t}^{\theta}$ as a time-condition function of a noisy microenvironment $\mathcal{M}$ and a source $\mathcal{M}^{0}$ with outputs:
@@ D.8 Structure-aware evaluation
While we refer to our primary evaluation metrics as PSD and SPD, they correspond to the two asymmetric directions of the Chamfer Distance (CD) - a widely used metric in point cloud generation [chamferdistance]. PSD measures how closely the predicted points adhere to the ground truth structure (fidelity), while SPD captures how well the prediction covers the full extent of the target (coverage). We compute both metrics over collections of microenvironments tiled across the tissue, providing an indirect assessment of global morphological accuracy.
To directly assess the preservation of internal structure within generated microenvironments, we additionally evaluate Gromov-Wasserstein (GW) and Fused Gromov-Wasserstein (FGW) distances [gw, vayer2020fused]. These metrics compare the intra-point pairwise distances in predicted and ground truth point clouds, capturing latent structural relationships. FGW further incorporates transcriptomic similarity, offering a holistic measure of structural and feature-based alignment.